OpenStudy (anonymous):

Maria looks at the architectural plan of a four-walled room in which the walls meet each other at right angles. The length of one wall in the plan is 13 inches. The length of the diagonal of the floor of the room in the plan is approximately 15.26 inches. Is the room in the shape of a square? Explain how you determined your answer. Show all your work.

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@YanaSidlinskiy @ganeshie8

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@.Sam.

OpenStudy (anonymous):

Hello, To find the diagonal length of a square, take a perimeter and multiply it by \(\sqrt{2}\). \[13 \times \sqrt{2} = 18.38477631085023563442195341472607502140573437990032495129683...\] \[18.38 \ne 15.26\] Hence, the room is not a square. Peace.

OpenStudy (anonymous):

o_O

OpenStudy (anonymous):

I want the answer in english please.

OpenStudy (anonymous):

Okay, let us start this off step by step. Do you know the formula for the diagonal length of a square?

OpenStudy (anonymous):

Ehh, no.

OpenStudy (anonymous):

Okay so learn it, the diagonal length of a square can be calculated by this formula: \[x \times \sqrt{2}\] Where \(x\) is one side of the perimeter. Do you follow and understand so far?

OpenStudy (anonymous):

Uhhh

OpenStudy (anonymous):

I think you're supposed to find the answer with the Pythagorean theorem...

OpenStudy (anonymous):

It asks one question, "Is the room in the shape of a square?" Using the formula, if the product does not equal the given diagonal length, the object is not a square.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I am basically saying, the use of the Pythagorean theorem here is unneeded, the question is simple enough to be solved with just a formula and basic arithmetic skills.

OpenStudy (anonymous):

Welp, that's what I'm supposed to use.

OpenStudy (anonymous):

Well then set up a right angle triangle (the question gives you the details) solve for the hypotenuse then compare it to the diagonal length they give you. If it matches it is square, if not, then it is not a square.